# ThermoSplat: Cross-Modal 3D Gaussian Splatting with Feature Modulation and Geometry Decoupling

### v1

@@ Abstract
Multi-modal scene reconstruction integrating RGB and thermal infrared data is essential for robust environmental perception across diverse lighting and weather conditions. However, extending 3D Gaussian Splatting (3DGS) to multi-spectral scenarios remains challenging. Current approaches often struggle to fully leverage the complementary information of multi-modal data, typically relying on mechanisms that either tend to neglect cross-modal correlations or leverage shared representations that fail to adaptively handle the complex structural correlations and physical discrepancies between spectrums. To address these limitations, we propose ThermoSplat, a novel framework that enables deep spectral-aware reconstruction through active feature modulation and adaptive geometry decoupling. First, we introduce a Cross-Modal FiLM Modulation mechanism that dynamically conditions shared latent features on thermal structural priors, effectively guiding visible texture synthesis with reliable cross-modal geometric cues. Second, to accommodate modality-specific geometric inconsistencies, we propose a Modality-Adaptive Geometric Decoupling scheme that learns independent opacity offsets and executes an independent rasterization pass for the thermal branch. Additionally, a hybrid rendering pipeline is employed to integrate explicit Spherical Harmonics with implicit neural decoding, ensuring both semantic consistency and high-frequency detail preservation. Extensive experiments on the RGBT-Scenes dataset demonstrate that ThermoSplat achieves state-of-the-art rendering quality across both visible and thermal spectrums.
@@ I Introduction
To bridge these gaps, we present ThermoSplat, a novel cross-modal 3DGS framework that enables deep spectral-aware reconstruction through active feature modulation and adaptive geometry decoupling. Unlike existing methods that either rely on explicit decomposition or unified latent representation, ThermoSplat introduces a FiLM-based [19] feature modulation mechanism. This module dynamically conditions the shared latent representation on thermal structural priors, enabling the model to actively leverage structural infrared features to guide visible texture synthesis. Furthermore, to accommodate the inherent physical discrepancies across different spectral bands, we propose a modality-adaptive geometric decoupling scheme, which allows for independent geometric adjustments in the infrared spectrum, effectively resolving artifacts in regions where transparency or reflectivity varies. Finally, to overcome the detail-loss inherent in pure feature-based decoding, we employ a hybrid rendering pipeline, which integrates explicit Spherical Harmonics (SH) with implicit decoding, achieving high-frequency RGB details while maintaining consistent semantic information across modalities.
@@
  • Cross-Modal FiLM Modulation: We design a Feature-wise Linear Modulation (FiLM) framework to establish deep feature dependencies. By utilizing structural priors to modulate shared latent features, our method enhances texture recovery and cross-modal alignment.
@@ III-A Overview
The overall architecture of ThermoSplat is designed to achieve high-fidelity multi-modal scene reconstruction by addressing spectral-varying properties. As illustrated in Fig. 1, we represent the scene using multi-modal feature-enhanced 3DGS [9]. The pipeline first performs active feature interaction via a Cross-Modal FiLM Modulation on the rasterized latent representations, which utilizes thermal structural priors to guide visible texture synthesis. To account for geometric inconsistencies across spectrums, we introduce a modality-adaptive geometric decoupling scheme, which uses the learnable offset $\Delta_{t}\alpha$ and executes an independent rasterization pass to accommodate modality-specific geometries. Finally, a hybrid rendering strategy is employed to combine explicit Spherical Harmonics (SH) with implicit feature-decoded outputs for preserving high-frequency details and view-dependent effects in the visible spectrum.
@@
Figure 1: Overview of the proposed ThermoSplat framework. Given multi-spectral inputs, our method optimizes 3D Gaussian primitives with decoupled properties. (a) Cross-Modal FiLM Modulation dynamically conditions shared latent features on thermal structural priors to guide visible texture synthesis. (b) Modality-Adaptive Geometric Decoupling resolves geometric inconsistencies between visible and infrared spectrums. (c) The Hybrid Rendering pipeline integrates explicit Spherical Harmonics (SH) with implicit neural decoding, ensuring high-frequency detail preservation and cross-modal semantic consistency.
@@ III-C Cross-Modal Feature Modulation
To address the spectral gap between visible and infrared modalities, we propose a cross-modal modulation mechanism. Instead of treating visible and thermal signals as independent entities, our framework leverages structural priors inherent in the thermal spectrum to guide the synthesis of visible textures. As shown in Fig. 1, the proposed Cross-Modal FiLM [19] Modulation integrates feature extraction and conditioning in a unified neural architecture.
@@
Thermal-Guided FiLM Modulation. Distinct from directly applying MLP layers for different spectrums, we introduce a Thermal Prior Head $\Phi_{th}$ and a FiLM-based modulation [19] layer. Crucially, as $h_{th}=\Phi_{th}(h)$ is directly supervised by the thermal rendering task through the subsequent decoding stage (Eq. 6), it is naturally driven to distill structural priors that are most representative of the infrared domain. The feature $h_{th}$ is subsequently mapped to a set of modulation parameters $(\gamma,\beta)$ through a linear transformation:
$\gamma,\beta]=\text{Linear}_{film}(h_{th}).$ (4)
By treating the infrared information as a conditioning signal, we apply Feature-wise Linear Modulation (FiLM) to the shared representation $h$:
@@
where $\mathcal{C}_{implicit}^{rgb}$ provides the base color component for the subsequent hybrid rendering stage. Notably, the thermal-specific feature $h_{th}$ serves a dual purpose: it acts as the source for FiLM parameter generation while simultaneously being decoded into the infrared signal $\mathcal{C}^{thermal}$. This hierarchical modulation ensures that the synthesis of visible images is physically constrained by the cross-modal structural consistency.
@@ IV-C Ablation
In this paper, we present ThermoSplat, a novel cross-modal 3D Gaussian Splatting framework designed for high-fidelity RGBT scene reconstruction. To effectively bridge the gap between visible and thermal modalities, we introduce a Cross-Modal FiLM Modulation mechanism that leverages thermal structural priors to guide visible texture synthesis. Furthermore, to address the inherent geometric inconsistencies caused by disparate physical sensing properties, we propose a Modality-Adaptive Geometric Decoupling scheme, which enables the model to accurately represent independent spectral characteristics without compromising spatial alignment. Extensive experiments on the RGBT-Scenes dataset demonstrate that our approach achieves state-of-the-art performance in both rendering quality and structural accuracy. By integrating explicit geometric representations with implicit neural feature modulation, ThermoSplat provides a robust and efficient solution for multi-spectral scene understanding in visually degraded environments.

### v2

@@ Abstract
Multi-modal scene reconstruction integrating RGB and thermal infrared data is essential for robust environmental perception across diverse lighting and weather conditions. However, extending 3D Gaussian Splatting (3DGS) to multi-spectral scenarios remains challenging. Current approaches often struggle to fully leverage the complementary information of multi-modal data, typically relying on mechanisms that either tend to neglect cross-modal correlations or leverage shared representations that fail to adaptively handle the complex structural correlations and physical discrepancies between spectrums. To address these limitations, we propose ThermoSplat, a novel framework that enables deep spectral-aware reconstruction through active feature modulation and adaptive geometry decoupling. First, we introduce a Spectrum-Aware Adaptive Modulation that dynamically conditions shared latent features on thermal structural priors, effectively guiding visible texture synthesis with reliable cross-modal geometric cues. Second, to accommodate modality-specific geometric inconsistencies, we propose a Modality-Adaptive Geometric Decoupling scheme that learns independent opacity offsets and executes an independent rasterization pass for the thermal branch. Additionally, a hybrid rendering pipeline is employed to integrate explicit Spherical Harmonics with implicit neural decoding, ensuring both semantic consistency and high-frequency detail preservation. Extensive experiments on the RGBT-Scenes dataset demonstrate that ThermoSplat achieves state-of-the-art rendering quality across both visible and thermal spectrums.
@@ I Introduction
To bridge these gaps, we present ThermoSplat, a novel cross-modal 3DGS framework that enables deep spectral-aware reconstruction through active feature modulation and adaptive geometry decoupling. Unlike existing methods that either rely on explicit decomposition or unified latent representation, ThermoSplat introduces a Spectrum-Aware Adaptive Modulation. While drawing inspiration from the mathematical flexibility of linear modulation [19], this module dynamically conditions the shared latent representation on thermal structural priors, enabling the model to actively leverage structural infrared features to guide visible texture synthesis. Furthermore, to accommodate the inherent physical discrepancies across different spectral bands, we propose a modality-adaptive geometric decoupling scheme, which allows for independent geometric adjustments in the infrared spectrum, effectively resolving artifacts in regions where transparency or reflectivity varies. Finally, to overcome the detail-loss inherent in pure feature-based decoding, we employ a hybrid rendering pipeline, which integrates explicit Spherical Harmonics (SH) with implicit decoding, achieving high-frequency RGB details while maintaining consistent semantic information across modalities.
@@
  • Spectrum-Aware Adaptive Modulation: We design a Spectrum-Aware Adaptive Modulation framework to establish deep feature dependencies. By utilizing structural priors to modulate shared latent features, our method enhances texture recovery and cross-modal alignment.
@@ III-A Overview
The overall architecture of ThermoSplat is designed to achieve high-fidelity multi-modal scene reconstruction by addressing spectral-varying properties. As illustrated in Fig. 1, we represent the scene using multi-modal feature-enhanced 3DGS [9]. The pipeline first performs active feature interaction via a Spectrum-Aware Adaptive Modulation on the rasterized latent representations, which utilizes thermal structural priors to guide visible texture synthesis. To account for geometric inconsistencies across spectrums, we introduce a modality-adaptive geometric decoupling scheme, which uses the learnable offset $\Delta_{t}\alpha$ and executes an independent rasterization pass to accommodate modality-specific geometries. Finally, a hybrid rendering strategy is employed to combine explicit Spherical Harmonics (SH) with implicit feature-decoded outputs for preserving high-frequency details and view-dependent effects in the visible spectrum.
@@
Figure 1: Overview of the proposed ThermoSplat framework. Given multi-spectral inputs, our method optimizes 3D Gaussian primitives with decoupled properties. (a) Spectrum-Aware Adaptive Modulation dynamically conditions shared latent features on thermal structural priors to guide visible texture synthesis. (b) Modality-Adaptive Geometric Decoupling resolves geometric inconsistencies between visible and infrared spectrums. (c) The Hybrid Rendering pipeline integrates explicit Spherical Harmonics (SH) with implicit neural decoding, ensuring high-frequency detail preservation and cross-modal semantic consistency.
@@ III-C Cross-Modal Feature Modulation
To address the spectral gap between visible and infrared modalities, we propose a cross-modal modulation mechanism. Instead of treating visible and thermal signals as independent entities, our framework leverages structural priors inherent in the thermal spectrum to guide the synthesis of visible textures. As shown in Fig. 1, the proposed Spectrum-Aware Adaptive Modulation integrates feature extraction and conditioning in a unified neural architecture.
@@
Spectrum-Aware Adaptive Modulation. Distinct from directly applying MLP layers for different spectrums, we introduce a Thermal Prior Head $\Phi_{th}$ and a Spectrum-Aware Adaptive Modulation [19] layer. Unlike generic conditioning methods, our modulation parameters $(\gamma,\beta)$ are dynamically derived from a dedicated Thermal Prior Head $\Phi_{th}$. Crucially, as $h_{th}=\Phi_{th}(h)$ is directly supervised by the thermal rendering task through the subsequent decoding stage (Eq. 6), the modulation process is implicitly physics-aware, ensuring that the distilled structural priors from the infrared domain actively guide the visible feature synthesis. The feature $h_{th}$ is subsequently mapped to a set of modulation parameters $(\gamma,\beta)$ through a linear transformation:
$\gamma,\beta]=\Phi_{mod}(h_{th}).$ (4)
By treating the infrared information as a conditioning signal, we apply Spectrum-Aware Adaptive Modulation to the shared representation $h$:
@@
where $\mathcal{C}_{implicit}^{rgb}$ provides the base color component for the subsequent hybrid rendering stage. Notably, the thermal-specific feature $h_{th}$ serves a dual purpose: it acts as the source for the proposed modulation parameter generation while simultaneously being decoded into the infrared signal $\mathcal{C}^{thermal}$. This hierarchical modulation ensures that the synthesis of visible images is physically constrained by the cross-modal structural consistency.
@@ IV-C Ablation
In this paper, we present ThermoSplat, a novel cross-modal 3D Gaussian Splatting framework designed for high-fidelity RGBT scene reconstruction. To effectively bridge the gap between visible and thermal modalities, we introduce a Spectrum-Aware Adaptive Modulation mechanism that leverages thermal structural priors to guide visible texture synthesis. Furthermore, to address the inherent geometric inconsistencies caused by disparate physical sensing properties, we propose a Modality-Adaptive Geometric Decoupling scheme, which enables the model to accurately represent independent spectral characteristics without compromising spatial alignment. Extensive experiments on the RGBT-Scenes dataset demonstrate that our approach achieves state-of-the-art performance in both rendering quality and structural accuracy. By integrating explicit geometric representations with implicit neural feature modulation, ThermoSplat provides a robust and efficient solution for multi-spectral scene understanding in visually degraded environments.